OpenStudy (anonymous):

Determine the prime factorization of 60. A. 22 • 11 B. 2 • 32 • 5 C. 23 • 11 D. 22 • 3 • 5

OpenStudy (anonymous):

@200205650

OpenStudy (danjs):

those 23 32 22 are really multiplied right?

OpenStudy (anonymous):

ohhh no there squared sorry forgot to mention that

OpenStudy (danjs):

oh i c , ok

OpenStudy (danjs):

since 60 is even , start dividing by 2 as many times as you can

OpenStudy (danjs):

60 = 2 * 30

OpenStudy (anonymous):

60/2=10

OpenStudy (danjs):

60/2 = 30...

OpenStudy (anonymous):

sorry about that. :p

OpenStudy (danjs):

keep pulling out 2's until you cant (have a odd number) 60 = 2 * 30 = 2 * 2 * 15

OpenStudy (anonymous):

60

OpenStudy (danjs):

you want to break 60 down into a product of prime numbers...

OpenStudy (anonymous):

its B

OpenStudy (danjs):

60 = 2 * 30 = 2 * 2 * 15 = 2 * 2 * 3 * 5

OpenStudy (danjs):

2, 3, and 5 are all prime... 60 = 2^2 * 3 * 5

OpenStudy (danjs):

Or, Since they give you multiple choice, you can see if only 1 totals 60 A) 2^2 * 11 = 44 B) 2 * 3^2 * 5 = 90 C) 2^3 * 11 = 88 D) 60